The current stock price of the Down Under Clothing Company of Australia is A$50 per share and its expected rate of return over the coming year is 14%. The market risk premium in Australia is 8% per year and the risk-free interest rate 6% per year. According to the CAPM, what would happen to the stock's current price if the covariance of the stock with the market portfolio falls by half while everthing else remians the same?
Answer #1

According to CAPM

Return = Risk free + Beta *(Rm - Rf)

14% = 6% + Beta * 8% (as 8% is premium. So Rm-Rf = 8%)

This implies Beta = 1

Now, if stock's covariance with market falls by half, so does beta and so it is now 0.5

Expected return now will be Rf + Beta * (Premium)

6% + 0.5 * 8% = 10%

Now, Since expected return was 14%, stock price after 1 year was expected to be 50*1.14 = $57.

Now, since the stocks expected return is only 10% and assuming stock will still achieve its next year,s stock price, the current share of price should be equal to 57/1.1 = $51.8

So stock price will increase to $51.8